A Bibliography of Publications in Computer Networks: The International Journal of Distributed Informatique

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

15 April 2020 Version 1.07

Title word cross-reference

n [52].

-Philosophers [52].

1982 [103].

4th [206].

5th [207].

6th [208].

70-X.25 [107]. 7th [209].

8th [210].

9th [211].

Academy [163]. Access [1, 49, 60, 86, 102, 105, 108, 119, 126]. Access-Control [105]. Accessing [125]. Acknowledging [100]. Activities [75]. Adaptable [131]. Adaptive [45, 136]. Addressing [158, 197]. Advanced [141, 164]. Airborne [164]. Algorithm [48, 80, 100, 138, 160]. Algorithms [47, 136]. Alis [121]. Allocation [19, 95]. Alphanumeric [14]. American [126]. Analysis [2, 7, 10, 61, 91, 99, 102, 130, 138, 140, 166, 170, 201]. Analytical [119]. ANSI [54]. ANUNET [59]. Application [125, 172, 175]. Applications [132, 148, 178]. Applying [85]. Approach [21, 54, 66]. Arbiter [165]. Architectural [189]. Architecture [50, 66, 85, 98, 114, 127, 151, 154, 182, 184]. Area [88, 89, 110, 117, 130, 132, 137–139, 155, 163, 175, 178, 181, 182, 186, 187, 189, 198, 201]. ARPA [8, 68]. Arpanet [48]. Artificial [110]. Aspects [117]. Assignment [72]. Australian [59]. Authentication [150, 167]. Automated [32, 150]. Automation [74]. Availability [135].

Balancing [194]. Base [16]. Based [1, 26, 30, 34, 63, 69, 78, 94, 95, 99, 131, 155, 158, 170, 182, 188]. Basic [64, 74]. Behavior [75]. Belgian [21]. Benefit [10]. Between [112]. Bit [174]. Broadband [151]. Broadcast [47, 51, 120, 159, 160]. Buffer [19, 58]. Bus [86, 118, 149, 165, 172, 200, 202–204]. BX.25 [93].

Cable [105, 172, 174]. Call [152]. Cambridge [81, 87]. Campus [190]. Can [96]. Cancellation [193]. Carrier [49]. CBCR [188]. CCITT [134]. CD [193, 194, 203, 205]. Centralized [173]. Certification [93]. Chain [32]. Change [133]. Channel [188]. Channels [102, 105, 167]. Characteristics [181, 195]. Choice [18]. Circuit [60]. Circuit/Packet [60]. Circuits [65]. CLAN [163]. Clock [160]. Clustering [176]. Clusters [170]. Co [127]. Co-Operating [127]. Coax [162]. Cobol [125]. College [198]. Collision [86, 105, 203]. Collision-Free [86, 105]. Combined [196]. Common [53, 174]. Communication [17, 18, 24, 30, 31, 34, 36, 56, 61, 62, 66, 69, 79, 87, 99, 129, 138, 167, 200]. Communications [6, 11, 42, 58, 71, 76]. Communications-Oriented [71]. Comparing [83]. Comparison [43, 87, 137]. Comparisons [56]. Complete [94]. Computer [4, 7, 61, 65, 67, 72, 78, 85, 92, 94, 95, 106, 118, 131, 138, 161– 163, 165, 169, 172, 176, 180, 185, 199, 206–211]. Computer-Based [94, 131]. Computer-Communication [61].

Computerized [112]. Computers [157]. Concentrator [80]. Concept [34, 118, 123]. Concepts [134, 182]. Conference [206–211]. Configuration [159]. Conflict [170]. Conflict-Free [170]. Connected [46]. Connecting [156]. Connection [38, 63]. Considerations [189]. Contention [100, 186, 188]. Contextual [197]. Control [22, 52, 61, 83, 105, 144, 158, 173]. Controlled [136]. Controllers [191]. Cooperation [99]. Cooperative [138]. Copying [73]. Corporation [180]. Correctness [96]. Cost [10, 169]. Cost-Benefit [10]. Coupled [118]. Criteria [11]. Criterion [101]. CSIRONET [161]. CSMA [159, 179, 187, 193, 194, 203]. CSMA-CD [205]. CSMA-Type [179]. CSMA/CD [193, 194, 203]. CSU [198]. Current [96].

Data [5, 6, 11, 14, 16, 18, 42, 45, 53, 56, 58, 64, 83, 108, 115, 129, 130, 175, 205]. Data-Entry [14]. Data-Sharing [6]. Database [126, 127, 135]. Databases [57]. Datagram [13, 20]. Decentralized [77, 136]. Decisions [8]. Defense [115]. Definition [35, 37, 166]. Delay [166]. Delta [127]. Denmark [121]. Department [115]. Dependency [135]. Dependent [159]. Description [27, 36]. Design [2, 8, 40, 46, 67, 80, 90, 101, 128, 129, 155, 156, 178, 179, 183]. Destination [158]. Development [32, 161]. Developments [59]. Dialogue [193]. Digital [44, 76, 87]. Din [131]. Din-A-Mit [131]. Discrete [180]. Distributed [15, 16, 40, 41, 47, 52, 57, 73, 104, 112, 121, 131, 135, 144, 146, 161, 173, 184]. Document [121, 134]. Documents [112]. DoD [114]. Draft [134]. DTDMA [130].

Economic [56]. Effect [42]. Efficient [105, 194]. Electronic [75]. Elementary [40]. Emulation [179]. Encapsulation [70]. End [25]. End-To-End [25]. Endpoint [152]. Engineering [162, 198]. Entry [14].

Environment [31, 78, 149, 152, 172, 177, 196]. Ethernet [87, 191]. Evaluating [135]. Evaluation [3, 11, 28, 191]. Event [180]. Eventcounts [51]. Example [121]. Execution [23]. Executive [164]. Exercise [52]. Experimental [77, 137]. Experiments [5].

Facilities [62]. Facility [93, 179]. Facing [74]. Facsimile [18]. Factors [185]. Fading [58]. FARA [205]. FARA/CSMA [205]. FARA/CSMA-CD [205]. Fast [192]. Fault [120]. Fault-Tolerant [120]. Fiber [118, 175]. File [29, 72, 95, 146]. Filestore [123]. Finite [36]. Firmware [168]. Flow [61, 107]. Flows [65]. Formal [98]. Fortran [124]. Forward [101]. Fragmentation [39, 82]. Framework [127]. Free [19, 86, 105, 170]. Ftp [29]. Functions [59, 69]. Future [148].

Gamma [118]. Gamma-Net [118]. Gateway [107]. Gateway-Network [107]. General [167]. Generating [110]. Generators [110]. Generic [184]. Growth [162]. Guarantee [96].

Hardware [154]. Haul [156]. HDLC [28, 83, 158]. Heterogeneous [157]. Hierarchical [3]. Hierarchically [46]. High [14, 118, 152, 188]. High-Speed [152]. Hilton [208–210]. Homogeneous [6, 7]. Hotel [211]. Human [185]. Hybrid [186, 204, 205]. Hybrinet [204]. Hyperchannel [180].

Identity [112]. IIASA [107]. Implementation [15, 20, 33, 51, 154, 168, 183]. Implementations [97]. Implemented [27]. Implications [113]. Improved [84]. Informal [43]. Information [18, 94, 107, 120, 141, 142]. Inhouse [116]. Inn [206–210]. INS [141, 143].

Instrumentation [132]. Integrated [79, 151]. Integration [205]. Inter [39]. Inter-Network [39]. Interactive [10, 129]. Interchange [134]. Interconnect [177]. Interconnection [4, 64, 106, 152, 196, 197]. Interface [20, 53, 124]. Interfaces [122, 200]. International [103, 107]. Internet [68, 114]. Internetwork [70]. Interprocess [31]. Interworking [117]. Investigation [19]. ISO [54]. ISO-ANSI [54]. Issues [74, 135, 156, 199].

Job [136]. July [103].

KEIO [190]. Key [199]. Knapsack [95].

Laboratories [12]. Laboratory [5, 177]. LAN [196]. LAN-PDN [196]. Language [126]. Large [3, 172]. Large-Scale [172]. Layer [145]. Level [14, 168]. Library [121]. Line [121]. Linear [174]. Link [83, 158, 168]. Link-Control [83]. Links [58]. Literature [121]. Load [194]. Local [1, 49, 69, 85, 86, 88, 89, 100, 110, 117, 118, 130, 132, 137–139, 148, 149, 151– 153, 155, 156, 160, 161, 163, 165, 169, 173– 176, 178, 179, 181–183, 185– 189, 198, 199, 201, 206–211]. Local-Area [182]. Localization [121]. Long [156]. Loop [42, 45, 192]. Loop-Type [45]. Low [169, 202]. Low-Overhead [202].

M [187]. M-CSMA [187]. Machine [55]. machines [1]. Mail [78]. Management [38, 63, 126]. Manufacturing [162]. Marquette [206, 207]. Measurements [22]. Measuring [75]. Mechanisms [26, 63, 145]. Media [17, 18]. Medium [169]. Memo [150]. Message [61, 69, 90, 111, 112, 131, 182, 183, 193]. Message-Based [69]. Message-Passing [182]. Message-Switched [111]. Method [95]. Microcomputers [155]. Micronet [154]. Microprocessor [33, 149]. Military [113, 148, 163, 178]. Mininet [132]. Minneapolis [206–211]. Minnesota [206–211]. MITS [33]. Mixed [17]. Mixed-Media [17]. Mobile [108]. Model [54, 64, 94, 114, 143, 180]. Modeling [61, 85, 99]. Modelling [25]. Models [99]. Modem [153]. Modes [83]. Modest [177]. Modularity [128]. Monitoring [173]. Movable [195]. Multi [172]. Multi-Computer [172]. Multiaccess [69]. Multichannel [187]. Multicomputer [90]. Multidestination [147]. Multigraduated [62]. Multilevel [61, 80]. Multiple [49, 102, 119, 200]. Multipoint [158]. Multiprocess [15]. Mutual [70].

Naming [197]. National [5, 59, 121, 126]. Net [99, 118]. Network [1, 5– 9, 20, 21, 27, 31, 39, 41, 45, 46, 50, 51, 55, 59, 67, 72, 85, 86, 88, 90, 92, 95, 106, 107, 110, 118, 130, 132, 141, 142, 146, 149, 151, 155, 157, 159, 161– 163, 165, 168, 169, 171–173, 176, 178, 180– 183, 190, 198, 204]. Networks [3, 4, 13, 17, 49, 53, 60, 61, 65, 80, 89, 91, 100, 101, 108, 113, 117, 137– 139, 148, 152, 153, 156, 160, 174, 175, 179, 185– 189, 192–194, 199–202, 206–211]. No [153]. No-Modem [153]. Node [170]. Nomographs [2]. Nonpersistent [187]. Notation [98]. Number [110].

Objects [73]. October [206–211]. Office [74–76, 79, 149, 150]. On-Line [121]. One [32]. Open [54, 64, 66, 122, 128]. Operating [50, 71, 127, 144, 182]. Operational [83]. Optic [175]. Optical [118, 149]. Optimal [95]. Optimization [3, 17, 67]. Ordering [104, 121]. Organizational [75, 133]. Oriented [71]. OSI [125]. Overhead [202]. Overheads [82]. Overview [44, 78].

Packet [12, 13, 24, 39, 46, 60, 91, 102, 106, 108, 113, 119, 152]. Packet-Switched [13, 113]. Packet-Switching [24, 46, 106]. Pandora [129]. Parallel [57]. Passing [182, 183, 202]. PBX [76]. PDN [196]. Performance [3, 28, 81, 83, 87, 92, 103, 130, 137, 138, 159, 179, 181, 192, 201]. Personal [176, 199]. Perspective [185]. Petri [99]. Philips [12]. Philosophers [52]. Physical [5]. Piggybacking [91]. Plaza [211]. Practical [117]. Preserving [112]. Principles [66]. Prioritized [105, 159, 193]. Priority [69]. Privacy [150]. Private [106]. Problem [95]. Problems [67]. Procedure [126, 147]. Process [30]. Processing [64]. Processor [1, 90, 177]. Processors [41]. Program [115]. Programmable [55]. Programming [122]. Programs [125]. Promotes [107]. Proposal [54]. Proposed [126]. Protection [111, 144]. Protocol [14, 22, 24, 26, 28, 29, 32, 34, 35, 42, 48, 49, 63, 68, 84, 96–98, 105, 113, 115, 134, 140, 145, 156– 158, 166, 168, 170, 171, 183, 188, 194, 205]. Protocols [23, 25, 27, 36, 38, 39, 43, 60, 70, 79, 89, 99, 103, 129, 167, 186, 187]. Prototype [154]. Provision [193]. Public [13, 53, 116].

Radio [58, 100, 102, 105, 108]. Random [102, 110]. Rate [175]. Rates [174]. Real [132, 138]. Real-Time [132, 138]. Recommendation [134]. Reference [54, 64, 94]. References [112]. Reliable [15, 26, 63]. Repeat [147]. Requests [104]. Requirements [113]. Research [12]. Reservation [188]. Resolution [186, 203]. Resource [144, 161]. Response [42, 86]. Results [83]. Ring [81, 87, 109, 118, 137, 140, 155, 157, 168, 181, 201, 204]. Ring-Based [155]. RNS [169]. Routing [3, 48, 120, 197].

S [163]. S&Tnet [190]. S.a [134]. Satellite [119]. Scale [172]. Scheduling [136]. Scheme [202]. Scientific [107]. Searching [57]. Security [145, 184]. Selecting [199]. Selection [62]. Selective [147]. Sense [49]. Servers [104]. Service [13, 124, 196]. Services [11, 50, 56, 125, 134]. SGB [27].

Sharing [6, 92, 161]. Shortest [166]. Shortest-Delay [166]. Signatures [44]. Siliconizing [88]. Simulating [172]. Simulation [5, 7, 19, 41, 46, 47, 85, 90, 98, 159]. Simulations [136]. Sirius [127]. Sirius-Delta [127]. Size [58]. Slot [195]. Society [141]. Solution [95]. Some [189, 195]. Sorting [146]. Source [158]. Source-Destination [158]. Specification [98]. Speech [139]. Speed [118, 152, 169, 188]. Squire [71]. Stability [102]. Standard [126]. Standardisation [128]. Standardization [103, 113, 115]. State [36, 98, 135]. Station [33]. Statistics [9, 92]. Status [103]. Store [101]. Store-And-Forward [101]. Strategies [108]. Structure [59, 118, 165]. Structured [73]. Structures [42]. Structuring [66]. Studies [179]. Study [46, 65, 81, 119, 159, 192, 200]. Suitable [100]. Support [50, 162]. Supporting [139]. Survey [16, 35]. Switched [13, 61, 111, 113]. Switching [12, 24, 46, 60, 102, 106]. Symbolic [23]. Symmetric [138]. Synchronisation [111]. Synchronization [47, 160]. System [50, 71, 73, 77, 85, 97, 102, 104, 111, 118, 121, 122, 129, 131, 141–143, 149, 182]. Systematic [66]. Systems [10, 15, 40, 54, 64, 66, 69, 79, 94, 98, 112, 116, 126–128, 144, 180, 184].

Tariff [62]. TDM [195]. Technique [84, 98]. Techniques [35, 46, 96]. Technological [121]. Technologies [142]. Technology [16]. Telecommunications [141]. Telecommuting [133]. Telematic [134]. Teletex [196]. Teletex-Service [196]. Terminal [21, 34, 37, 125, 177]. Terminal-Processor [177]. Terminals [14]. Terranet [177]. Testing [97, 98]. Their [103, 113]. Theoretical [137]. Three [43, 136]. Time [42, 92, 132, 138]. Time-Sharing [92]. Timed [181]. Timer [26, 63]. Timer-Based [26, 63]. Timing [171]. Token [137, 140, 181, 201, 202, 204]. Token-Passing [202]. Tolerant [120]. TPA [107]. TPA/70 [107]. TPA/70-X.25 [107]. Traffic [110, 139]. Transfer [29]. Transmission [18, 22, 140, 175]. Transnetwork [82]. Transport [25, 33, 38, 63, 103, 124, 145]. Two [191]. Type [45, 179].

U. [163]. Unibus [191]. Unified [85]. University [21, 59]. UNIX [1, 190]. Unix-based [1]. Update [160]. Updating [48]. Upon [34]. Usage [9]. User [122]. Users [196]. Using [2, 23, 41, 46, 47, 99, 110].

Validation [32, 84, 180]. Variable [34, 175]. Vendor [106]. Verification [23, 25, 35, 96, 147]. Versus [116]. Videotex [116]. Violet [77]. Virtual [21, 34, 37, 55, 65, 123, 125, 152]. Voice [24, 130, 140, 205]. Voice-Data [205].

Workstation [176].

X.25 [53, 107, 124]. X25 [20, 30, 31, 89].

References

Manning:1976:UBL

[1] E. G. Manning, R. Howard, C. G. O’Donnell, K. Pammett, and E. Chang. A Unix-based local processor and network access machines. Computer Networks: The International Journal of Distributed Informatique, 1(2):139–142, September 1976. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Thananitayaudom:1977:ADU

[2] Tavorn Thananitayaudom. Analysis and design using nomographs. Computer Networks: The International Journal of Distributed Informatique, 1(3):147– 154, January 1977. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Kleinrock:1977:HRL

[3] Leonard Kleinrock and Fraouk Kamoun. Hierarchical routing for large networks. performance evaluation and optimization. Computer Networks: The International Journal of Distributed Informatique, 1(3):155–174, January 1977. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Sunchine:1977:ICN

[4] Carl A. Sunchine. Interconnection of computer networks. Computer Networks: The International Journal of Distributed Informatique, 1(3):175–195, January 1977. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Price:1977:DNS

[5] Wyn L. Price. Data network simulation. experiments at the National Physical Laboratory 1968–76. Computer Networks: The International Journal of Distributed Informatique, 1(4):199– 210, May 1977. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Manning:1977:HND

[6] Eric G. Manning and Richard W. Peebles. Homogeneous network for datasharing communications. Computer Networks: The International Journal of Distributed Informatique, 1(4):211– 224, May 1977. CODEN CNETDP.

ISSN 0376-5075 (print), 1878-3120 (electronic).

Labetoulle:1977:HCN

[7] Jacques Labetoulle, Eric G. Manning, and Richard W. Peebles. Homogeneous computer network. analysis and simulation. Computer Networks: The International Journal of Distributed Informatique, 1(4):225–240, May 1977. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

McQuillan:1977:AND

[8] John M. McQuillan and David C. Walden. ARPA network design decisions. Computer Networks: The International Journal of Distributed Informatique, 1(5):243–289, August 1977. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Lance:1977:NUS

[9] Godfrey N. Lance and A. Clive Edington. Network usage statistics. Computer Networks: The International Journal of Distributed Informatique, 1(6):295–305, November 1977. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Cotton:1977:CAI

[10] Ira W. Cotton. Cost-benefit analysis of interactive systems. Computer Networks: The International Journal of Distributed Informatique, 1(6):311–324, November 1977. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Grubb:1977:CED

[11] Dana S. Grubb and Ira W. Cotton. Criteria for evaluation of data communications services. Computer Networks: The International Journal of Distributed Informatique, 1(6):325–340, November 1977. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Burnett:1977:PSP

[12] D. J. Burnett and H. R. Sethi. Packet switching at Philips Research Laboratories. Computer Networks: The International Journal of Distributed Informatique, 1(6):341–348, November 1977. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Manning:1978:DSP

[13] Eric G. Manning. On datagram service in public packet-switched networks. Computer Networks: The International Journal of Distributed Informatique, 2(2):79–83, May 1978. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Naffah:1978:HLP

[14] Najah Naffah. High level protocol for alphanumeric data-entry terminals. Computer Networks: The International Journal of Distributed Informatique, 2(2):84–94, May 1978. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Lamport:1978:IRD

[15] Leslie Lamport. Implementation of reliable distributed multiprocess systems. Computer Networks: The International Journal of Distributed Informatique, 2(2):95–114, May 1978. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Davenport:1978:DDB

[16] Robert A. Davenport. Distributed data base technology — a survey. Computer Networks: The International Journal of Distributed Informatique, 2(3):155– 167, July 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Maruyama:1978:OMC

[17] Kiyoshi Maruyama. Optimization of mixed-media communication networks. Computer Networks: The International Journal of Distributed Informatique, 2 (3):168–178, July 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Kirstein:1978:CDC

[18] Peter T. Kirstein. Choice of data communication media for transmission of facsimile information. Computer Networks: The International Journal of Distributed Informatique, 2(3):179– 190, July 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Giessler:1978:FBA

[19] A. Giessler, J. Haenle, A. Koenig, and E. Pade. Free buffer allocation — an investigation by simulation. Computer Networks: The International Journal of Distributed Informatique, 2(3):191– 208, July 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Barber:1978:IXI

[20] Derek Barber, L. A., Tomaz Kalin, and C. M. Solomonides. Implementation of the X25 interface in a datagram network. Computer Networks: The International Journal of Distributed Informatique, 2(4–5):??, September/October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Bauwens:1978:VTA

[21] Eric G. Bauwens and Francis N. Magnee. Virtual terminal approach in the Belgian University Network. Computer Networks: The International Journal of Distributed Informatique, 2(4–5):??, September/October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Bennett:1978:MTC

[22] Christopher J. Bennett and Andrew J. Hinchley. Measurements of the Transmission Control Protocol. Computer Networks: The International Journal of Distributed Informatique, 2(4–5):??, September/October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Brand:1978:VPU

[23] Daniel Brand and William H. Joyner, Jr. Verification of protocols using symbolic execution. Computer Networks: The International Journal of Distributed Informatique, 2(4–5):??, September/October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Cohen:1978:PPV

[24] Dan Cohen. Protocol for packetswitching voice communication. Computer Networks: The International Journal of Distributed Informatique, 2 (4–5):??, September/October 1978. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Danthine:1978:MVE

[25] Andre Danthine and Joseph Bremer. Modelling and verification of end-to-end transport protocols. Computer Networks: The International Journal of Distributed Informatique, 2(4–5):??, September/October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Fletcher:1978:MRT

[26] John G. Fletcher and Richard W. Watson. Mechanisms for a reliable timerbased protocol. Computer Networks: The International Journal of Distributed Informatique, 2(4–5):??, September/ October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Gaudissart:1978:DPI

[27] E. M. Gaudissart, Stephane de Hepcee, Guy Melaet, and Vincent Obozinski. Description of protocols implemented in the SGB network. Computer Networks: The International Journal of Distributed Informatique, 2(4–5):??, September/October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Gelenbe:1978:PEH

[28] Erol Gelenbe, Jacques Labetoulle, and Guy Pujolle. Performance evaluation of the HDLC protocol. Computer Networks: The International Journal of Distributed Informatique, 2(4–5):??, September/October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Gien:1978:FTP

[29] Michel Gien. File Transfer Protocol (ftp). Computer Networks: The International Journal of Distributed Informatique, 2(4–5):??, September/October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Hertweck:1978:XBP

[30] Friedrich Hertweck, Eckart Raubold, and Fritz Vogt. X25 based process — process communication. Computer Networks: The International Journal of Distributed Informatique, 2(4–5):??, September/October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Jacquemart:1978:NIC

[31] Yves A. Jacquemart. Network interprocess communication in an X25 environment. Computer Networks: The International Journal of Distributed Informatique, 2(4–5):??, September/October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Rudin:1978:APV

[32] Harry Rudin, Colin H. West, and Pitro Zafiropulo. Automated protocol validation: One chain of development. Computer Networks: The International Journal of Distributed Informatique, 2 (4–5):??, September/October 1978. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Saettone:1978:MMI

[33] Robert Saettone. MITS: Microprocessor Implementation of a Transport Station. Computer Networks: The International Journal of Distributed Informatique, 2 (4–5):??, September/October 1978. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Schulze:1978:VTP

[34] Guenter Schulze and Joachim Boerger. Virtual terminal protocol based upon the “communication variable” concept. Computer Networks: The International Journal of Distributed Informatique, 2 (4–5):??, September/October 1978. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Sunshine:1978:SPD

[35] Carl A. Sunshine. Survey of protocol definition and verification techniques. Computer Networks: The International Journal of Distributed Informatique, 2 (4–5):??, September/October 1978. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Bochmann:1978:FSD

[36] Gregor v. Bochmann. Finite state description of communication protocols. Computer Networks: The International Journal of Distributed Informatique, 2(4–5):361–372, September/ October 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic). Also published in Welding in the World, Le Soudage Dans Le Monde, Feb 13–15 1978, Univ of Liège p. F3.1–F3.11.

Schicker:1978:VTD

[37] P. Schicker and A. Duenki. Virtual terminal definition. Computer Networks: The International Journal of Distributed Informatique, 2(6):429–441, December 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Sunshine:1978:CMT

[38] Carl A. Sunshine and Yogen K. Dalal. Connection management in transport protocols. Computer Networks: The International Journal of Distributed Informatique, 2(6):454–473, December 1978. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Shoch:1979:PFI

[39] John F. Shoch. Packet fragmentation in inter-network protocols. Computer Networks: The International Journal of Distributed Informatique, 3(1):3–8, February 1979. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Yeh:1979:DED

[40] Randolph T. Yeh and K. Mani Chandy. On the design of elementary distributed systems. Computer Networks: The International Journal of Distributed Informatique, 3(1):24–35, February 1979. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Peacock:1979:DSU

[41] J. Kent Peacock, J. W. Wong, and Eric G. Manning. Distributed simulation using a network of processors. Computer Networks: The International Journal of Distributed Informatique, 3 (1):44–56, February 1979. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Richardson:1979:EPR

[42] T. G. Richardson and Wilfred L. W. Yu. Effect of protocol on the response time of loop structures for data communications. Computer Networks: The International Journal of Distributed Informatique, 3(1):57–66, February 1979. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Postel:1979:ICT

[43] Jonathan B. Postel. Informal comparison of three protocols. Computer Networks: The International Journal of Distributed Informatique, 3(1):67–76, February 1979. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Matyas:1979:DSO

[44] Stephen M. Matyas. Digital signatures — an overview. Computer Networks: The International Journal of Distributed Informatique, 3(2):87–94, April 1979. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Yu:1979:ALD

[45] Wilfred L. W. Yu and Jayanti C. Majithia. Adaptive loop-type data network. Computer Networks: The International Journal of Distributed Informatique, 3(2):95–104, April 1979. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Gomberg:1979:DSH

[46] Gregory R. A. Gomberg, Ian H. Kerr, Julian S. Richards, Wyn L. Price, and C. M. Solomonides. Design study of a hierarchically connected packetswitching network using simulation techniques. Computer Networks: The International Journal of Distributed Informatique, 3(2):114–135, April 1979. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Peacock:1980:SDS

[47] J. Kent Peacock, Eric Manning, and J. W. Wong. Synchronization of distributed simulation using broadcast algorithms. Computer Networks: The International Journal of Distributed Informatique, 4(1):3–10, February 1980. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Rosen:1980:UPA

[48] Eric C. Rosen. Updating protocol of Arpanet’s new routing algorithm. Computer Networks: The International Journal of Distributed Informatique, 4 (1):11–19, February 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Lam:1980:CSM

[49] Simon S. Lam. Carrier Sense Multiple Access Protocol for local networks. Computer Networks: The International Journal of Distributed Informatique, 4 (1):21–32, February 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Watson:1980:ASN

[50] Richard W. Watson and John G. Fletcher. Architecture for support of network operating system services. Computer Networks: The International Journal of Distributed Informatique, 4 (1):33–49, February 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Bullis:1980:IEB

[51] Kenneth Bullis and William Franta. Implementation of eventcounts in a broadcast network. Computer Networks: The International Journal of Distributed Informatique, 4(2):57–69, April 1980. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Chang:1980:NED

[52] Ernest Chang. n-philosophers: An exercise in distributed control. Computer Networks: The International Journal of Distributed Informatique, 4(2):71– 76, April 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Rybczynski:1980:CXI

[53] Antony M. Rybczynski and John D. Palframan. Common X.25 interface to public data networks. Computer Networks: The International Journal of Distributed Informatique, 4(3):97–110, June 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Platowski:1980:IOS

[54] Thomas F. Platowski. ISO-ANSI Open Systems Reference Model — a proposal for a systems approach. Computer Networks: The International Journal of Distributed Informatique, 4(3):111– 124, June 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Kent:1980:PNV

[55] S. A. Kent. Programmable network virtual machine. Computer Networks: The International Journal of Distributed Informatique, 4(3):125–137, June 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Ratz:1980:ECD

[56] Herbert C. Ratz and James A. Field. Economic comparisons of data communication services. Computer Networks: The International Journal of Distributed Informatique, 4(4):143–155, September 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Srinivasan:1980:PSD

[57] B. Srinivasan. Parallel searching in distributed databases. Computer Networks: The International Journal of Distributed Informatique, 4(4):157–166, September 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Cavers:1980:BSD

[58] James K. Cavers and C. M. Woodside. Buffer size for data communications over fading radio links. Computer Networks: The International Journal of Distributed Informatique, 4(4):167–174, September 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Wolfendale:1980:AAN

[59] Garth L. Wolfendale. ANUNET — the Australian National University Network: Structure, functions and developments. Computer Networks: The International Journal of Distributed Informatique, 4(6):265–269, December 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Yu:1980:APC

[60] Wilfred Yu, Jayanti C. Majithia, and J. W. Wong. Access protocols for circuit/packet switching networks. Computer Networks: The International Journal of Distributed Informatique, 4 (6):271–283, December 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Georganas:1980:MAM

[61] Nicolas D. Georganas. Modeling and analysis of message switched computercommunication networks with multilevel flow control. Computer Networks: The International Journal of Distributed Informatique, 4(6):285–294, December 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Minoli:1980:SCF

[62] Daniel Minoli. Selection of communication facilities under a multigraduated tariff. Computer Networks: The International Journal of Distributed Informatique, 4(6):295–301, December 1980. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Watson:1981:TMR

[63] Richard W. Watson. Timer-based mechanisms in reliable transport protocol connection management. Computer Networks: The International Journal of Distributed Informatique, 5(1):47–56, February 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Anonymous:1981:DPO

[64] Anonymous. Data processing — Open Systems Interconnection — basic reference model. Computer Networks: The International Journal of Distributed Informatique, 5(2):81–118, April 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Labetoulle:1981:SFT

[65] Jacques Labetoulle and Guy Pujolle. Study of flows through virtual circuits computer networks. Computer Networks: The International Journal of Distributed Informatique, 5(2):119– 126, April 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Burkhardt:1981:SPC

[66] H. J. Burkhardt and S. Schindler. Structuring principles of the communication architecture of open systems — systematic approach. Computer Networks: The International Journal of Distributed Informatique, 5(3):157–166, May 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Samoylenko:1981:OPC

[67] S. I. Samoylenko. Optimization problems in computer network design. Computer Networks: The International Journal of Distributed Informatique, 5 (3):177–186, May 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Postel:1981:AIP

[68] Jonthan B. Postel, Carl A. Sunshine, and Danny Cohen. ARPA Internet Protocol. Computer Networks: The International Journal of Distributed Informatique, 5(4):261–271, July 1981. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Rom:1981:MPF

[69] Raphael Rom and Fouad A. Tobagi. Message-based priority functions in local multiaccess communication systems.

Computer Networks: The International Journal of Distributed Informatique, 5 (4):273–286, July 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Shoch:1981:MEI

[70] John F. Shoch, Danny Cohen, and Edward A. Taft. Mutual encapsulation of internetwork protocols. Computer Networks: The International Journal of Distributed Informatique, 5(4):287– 301, July 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Chesley:1981:SCO

[71] Harry R. Chesley and Bruce V. Hunt. Squire — a communications-oriented operating system. Computer Networks: The International Journal of Distributed Informatique, 5(5):333–339, September 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Foster:1981:FAC

[72] Derrell V. Foster, Lawrence W. Dowdy, and James E. Iv Ames. File assignment in a computer network. Computer Networks: The International Journal of Distributed Informatique, 5(5):341–349, September 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Sollins:1981:CSO

[73] Karen R. Sollins. Copying structured objects in a distributed system. Computer Networks: The International Journal of Distributed Informatique, 5 (5):351–359, September 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Panko:1981:FBI

[74] Raymond R. Panko. Facing basic issues in office automation. Computer Networks: The International Journal of Distributed Informatique, 5(6):391–399, December 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Conrath:1981:EOO

[75] David W. Conrath, Chris Higgins, Cherian S. Thachenkary, and William Wright. Electronic office and organizational behavior — measuring office activities. Computer Networks: The International Journal of Distributed Informatique, 5(6):401–410, December 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Richer:1981:OCD

[76] Ira Richer and Marianne Steiner. Office communications and the digital PBX. Computer Networks: The International Journal of Distributed Informatique, 5 (6):411–422, December 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Gifford:1981:VED

[77] David K. Gifford. Violet, an experimental decentralized system. Computer Networks: The International Journal of Distributed Informatique, 5(6):423–433, December 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Schicker:1981:CBM

[78] Peter Schicker. Computer based mail environment — an overview. Computer Networks: The International Journal of Distributed Informatique, 5(6):435–443, December 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Naffah:1981:CPI

[79] Najah Naffah. Communication protocols for integrated office systems. Computer Networks: The International Journal of Distributed Informatique, 5(6):445–454, December 1981. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Schneider:1982:ADM

[80] G. Michael Schneider and Mary N. Zastrow. Algorithm for the design of multilevel concentrator networks. Computer Networks: The International Journal of Distributed Informatique, 6(1):1–11, February 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Blair:1982:PSC

[81] Gordon S. Blair. Performance study of the Cambridge Ring. Computer Networks: The International Journal of Distributed Informatique, 6(1):13–20, February 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Bennett:1982:OTF

[82] Chris J. Bennett. Overheads of transnetwork fragmentation. Computer Networks: The International Journal of Distributed Informatique, 6(1):21–36, February 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Bux:1982:DLP

[83] Werner Bux, Karl Kuemmerle, and Hong Linh Truong. Data link-control performance: Results comparing HDLC operational modes. Computer Networks: The International Journal of Distributed Informatique, 6(1):37–51, February 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Rubin:1982:IPV

[84] J. Rubin and Colin H. West. Improved protocol validation technique. Computer Networks: The International Journal of Distributed Informatique, 6(2):65– 73, May 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Didic:1982:SLC

[85] Milena Didic and Bernd Wolfinger. Simulation of a local computer network architecture applying a unified modeling system. Computer Networks: The International Journal of Distributed Informatique, 6(2):75–91, May 1982. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Hamacher:1982:ARC

[86] V. Carl Hamacher and Gerald S. Shedler. Access response on a collisionfree local bus network. Computer Networks: The International Journal of Distributed Informatique, 6(2):93– 103, May 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Blair:1982:PCE

[87] Gordon S. Blair and Doug Shepherd. Performance comparison of Ethernet and the Cambridge Digital Communication Ring. Computer Networks: The International Journal of Distributed Informatique, 6(2):105–113, May 1982. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Coleman:1982:SLA

[88] Vernon Coleman. Siliconizing the local area network. Computer Networks: The International Journal of Distributed Informatique, 6(4):245–254, September 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Grant:1982:XPL

[89] Alastair Grant and David Hutchison. X25 protocols and local area networks. Computer Networks: The International Journal of Distributed Informatique, 6 (4):255–262, September 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Unger:1982:DSM

[90] Brian W. Unger and Don S. Bidulock. Design and simulation of a multicomputer network message processor. Computer Networks: The International Journal of Distributed Informatique, 6 (4):263–277, September 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Lai:1982:APP

[91] Wai Sum Lai. Analysis of piggybacking in packet networks. Computer Networks: The International Journal of Distributed Informatique, 6(4):279–290, September 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Iyengar:1982:PST

[92] S. Sitharama Iyengar and Wendy Chih Chun Liu. Performance statistics of a time-sharing computer network. Computer Networks: The International Journal of Distributed Informatique, 6(5):303–317, November 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Melici:1982:BCF

[93] Joseph A. Melici. BX.25 certification facility. Computer Networks: The International Journal of Distributed Informatique, 6(5):319–329, November 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Bachman:1982:TMC

[94] Charles W. Bachman and Ronald G. Ross. Toward a more complete reference model of computer-based information systems. Computer Networks: The International Journal of Distributed Informatique, 6(5):331–343, November 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Ceri:1982:OFA

[95] Stefano Ceri, Giancarlo Martella, and Giuseppe Pelagatti. Optimal file allocation in a computer network: a solution method based on the knapsack problem. Computer Networks: The International Journal of Distributed Informatique, 6(5):345–357, November 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Vemini:1982:CCP

[96] Yechiam Vemini and James F. Kurose. Can current protocol verification techniques guarantee correctness? Computer Networks: The International Journal of Distributed Informatique, 6 (6):377–381, December 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Rayner:1982:STP

[97] David Rayner. System for testing protocol implementations. Computer Networks: The International Journal of Distributed Informatique, 6(6):383–395, December 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Piatkowski:1982:SAN

[98] Thomas F. Piatkowski, Lap-Kin Ip, and Dayun He. State architecture notation and simulation: a formal technique for the specification and testing of protocol systems. Computer Networks: The International Journal of Distributed Informatique, 6(6):397–418, December 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Diaz:1982:MAC

[99] Michel Diaz. Modeling and analysis of communication and cooperation protocols using ET Petri net based models. Computer Networks: The International Journal of Distributed Informatique, 6(6):419–441, December 1982. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Malcolm:1983:ACA

[100] Michael A. Malcolm, Lawrence D. Rogers, and John E. Spracklen. Acknowledging contention algorithm suitable for local radio networks. Computer Networks: The International Journal of Distributed Informatique, 7(1):1–8, February 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Minoli:1983:NDC

[101] Daniel Minoli. New design criterion for store-and-forward networks. Computer Networks: The International Journal of Distributed Informatique, 7(1):9–15, February 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Szpankowski:1983:PSM

[102] Wojciech Szpankowski. Packet switching in multiple radio channels: Analysis and stability of a random access system. Computer Networks: The International Journal of Distributed Informatique, 7(1):17–26, February 1983. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

vonStudnitz:1983:TPT

[103] Peter von Studnitz. Transport protocols — their performance and status in international standardization (July 1982). Computer Networks: The International Journal of Distributed Informatique, 7(1):27–35, February 1983. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Piquer:1983:ORS

[104] Alfredo R. Piquer. On ordering requests to servers in a distributed system. Computer Networks: The International Journal of Distributed Informatique, 7(2):69–81, April 1983. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Gold:1983:ECP

[105] Yaron I. Gold and William F. Franta. Efficient collision-free protocol for prioritized access-control of cable or radio channels. Computer Networks: The International Journal of Distributed Informatique, 7(2):83–98, April 1983. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Fernandez:1983:IPP

[106] Joseph I. Fernandez and John A. Gibbons. Interconnection of a private packet-switching network and a computer vendor’s network. Computer Networks: The International Journal of Distributed Informatique, 7(2):99– 111, April 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Labadi:1983:ITG

[107] A. Labadi and Istvan Sebestyen. IIASA TPA/70-X.25 gateway-network promotes international flow of scientific information. Computer Networks: The International Journal of Distributed Informatique, 7(2):113–121, April 1983. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Riordon:1983:ASP

[108] J. S. Riordon, S. A. Mahmoud, S. E. Aidarous, and M. Niktash. Access strategies in packet mobile radio data networks. Computer Networks: The International Journal of Distributed Informatique, 7(4):211–221, August 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Saltzer:1983:WR

[109] J. H. Saltzer, K. T. Pogran, and D. D.

Clark. Why a ring? Computer Networks: The International Journal of Distributed Informatique, 7(4):223– 231, August 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Ramshaw:1983:GAT

[110] Lance A. Ramshaw and Paul D. Amer.

Generating artificial traffic over a local area network using random number generators. Computer Networks: The International Journal of Distributed Informatique, 7(4):233–251, August 1983. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Livesey:1983:PSM

[111] Jon Livesey and Eric Manning. Protection and synchronisation in a messageswitched system. Computer Networks: The International Journal of Distributed Informatique, 7(4):253–268, August 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Palme:1983:PIR

[112] Jacob Palme. On preserving identity of the references between documents in distributed computerized message systems. Computer Networks: The International Journal of Distributed Informatique, 7(4):269–272, August 1983. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Cerf:1983:MRP

[113] Vinton G. Cerf and Robert E. Lyons.

Military requirements for packetswitched networks and their implications for protocol standardization. Computer Networks: The International Journal of Distributed Informatique, 7 (5):293–306, October 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Cerf:1983:DIA

[114] Vinton G. Cerf and Edward A. Cain. The DoD Internet architecture model. Computer Networks: The International Journal of Distributed Informatique, 7 (5):307–318, October 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Selvaggi:1983:DDD

[115] Philip S. Selvaggi. Department of Defense Data Protocol Standardization Program. Computer Networks: The International Journal of Distributed Informatique, 7(5):319–328, October 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Maurer:1983:IVP

[116] H. A. Maurer and I. Sebestyen. Inhouse versus public videotex systems. Computer Networks: The International Journal of Distributed Informatique, 7 (5):329–342, October 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Heard:1983:LAN

[117] Kenneth S. Heard. Local area networks and the practical aspects of interworking. Computer Networks: The International Journal of Distributed Informatique, 7(5):343–348, October 1983. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Ebihara:1983:GLC

[118] Yoshihiko Ebihara, Katsuo Ikeda, Tomoo Nakamura, Ishizaka, Michihiro, Makoto Shinzawa, and Kazuhiko Nakayama. Gamma-net: a local computer network coupled by a high speed optical fiber ring bus — system concept and structure. Computer Networks: The International Journal of Distributed Informatique, 7(6):375–388, December 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Guihur:1983:ASP

[119] R. Guihur and J. Pellaumail. Analytical study of packet satellite multiple access. Computer Networks: The International Journal of Distributed Informatique, 7(6):389–393, December 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Perlman:1983:FBR

[120] Radia Perlman. Fault-tolerant broadcast of routing information. Computer Networks: The International Journal of Distributed Informatique, 7(6):395–405, December 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Barnholdt:1983:AOL

[121] B. Barnholdt and Jorgen Nybroe. Alis — the on-line literature localization and document ordering system of the National Technological Library of Denmark. an example of a distributed system. Computer Networks: The International Journal of Distributed Informatique, 7(6):407–411, December 1983. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Langsford:1984:OSU

[122] A. Langsford. Open system user’s programming interfaces. Computer Networks: The International Journal of Distributed Informatique, 8(1):3–12, February 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Linington:1984:VFC

[123] Peter F. Linington. Virtual filestore concept. Computer Networks: The International Journal of Distributed Informatique, 8(1):13–16, February 1984. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Andreoni:1984:FIX

[124] Gaetano Andreoni. Fortran interface to X.25 and transport service. Computer Networks: The International Journal of Distributed Informatique, 8(1):17–22, February 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Dhondy:1984:AOV

[125] Rohinton Dhondy, Michel Elie, and Francis Touzin. Accessing OSI virtual terminal services from Cobol application programs. Computer Networks: The International Journal of Distributed Informatique, 8(1):23–29, February 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Gallagher:1984:PLA

[126] Leonard J. Gallagher. Procedure language access to Proposed American National Standard Database Management Systems. Computer Networks: The International Journal of Distributed Informatique, 8(1):31–42, February 1984.

CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Esculier:1984:SAF

[127] Christian Esculier. Sirius-Delta architecture: a framework for co-operating database systems. Computer Networks: The International Journal of Distributed Informatique, 8(1):43–48, February 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Beech:1984:MDS

[128] David Beech. Modularity in the design and standardisation of open systems. Computer Networks: The International Journal of Distributed Informatique, 8(1):49–55, February 1984. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Holzmann:1984:PSI

[129] Gerard J. Holzmann. Pandora System: an interactive system for the design of data communication protocols. Computer Networks: The International Journal of Distributed Informatique, 8(2):71–79, April 1984. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Li:1984:PAD

[130] San qi Li and J. C. Majithia. Performance analysis of a DTDMA local area network for voice and data. Computer Networks: The International Journal of Distributed Informatique, 8(2):81– 91, April 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Bauerfeld:1984:DAD

[131] Wulf L. Bauerfeld, Peter Kaufmann, Michael Tschicholz, and Chris Weikart. Din-A-Mit: an adaptable, distributed computer-based message system. Computer Networks: The International Journal of Distributed Informatique, 8(2):93–105, April 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Neri:1984:MLA

[132] Giovanni Neri, Richard C. S. Morling, Gerald D. Cain, Eugenio Faldella, Marco Longhi-Gelati, Tullio Salmon-Cinotti, and Pierantonio Natali. Mininet: a local area network for real-time instrumentation applications. Computer Networks: The International Journal of Distributed Informatique, 8(2):107–131, April 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Farwell:1984:TOC

[133] Barbara M. Farwell and David C. Farwell. Telecommuting and organizational change. Computer Networks: The International Journal of Distributed Informatique, 8(3):169–173, June 1984. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Horak:1984:CDI

[134] Wolfgang Horak. Concepts of the Document Interchange Protocol for the Telematic Services — CCITT draft recommendation S.a. Computer Networks: The International Journal of Distributed Informatique, 8(3):175–185, June 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Schreiber:1984:SDI

[135] Fabio A. Schreiber. State dependency issues in evaluating distributed database availability. Computer Networks: The International Journal of Distributed Informatique, 8(3):187–197, June 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Stankovic:1984:STA

[136] John A. Stankovic. Simulations of three adaptive, decentralized controlled, job scheduling algorithms. Computer Networks: The International Journal of Distributed Informatique, 8(3):199– 217, June 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Sventek:1984:TRL

[137] J. Sventek, W. Greiman, M. O’Dell, and A. Jansen. Token ring local area networks: a comparison of experimental and theoretical performance. Computer Networks: The International Journal of Distributed Informatique, 8(4):301– 309, August 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Mahbod:1984:PAC

[138] B. Mahbod and J. Howard. Performance analysis of a cooperative symmetric algorithm for real-time local area computer communication networks. Computer Networks: The International Journal of Distributed Informatique, 8 (4):311–323, August 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Hilal:1984:LAN

[139] Wael Hilal and Ming T. Liu. Local area networks supporting speech traffic. Computer Networks: The International Journal of Distributed Informatique, 8(4):325–337, August 1984. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Wong:1984:ATR

[140] J. W. Wong and P. M. Gopal. Analysis of a token ring protocol for voice transmission. Computer Networks: The International Journal of Distributed Informatique, 8(4):339–346, August 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Kitahara:1984:IIN

[141] Yasusada Kitahara. INS (Information Network System) — telecommunications for the advanced information society. Computer Networks: The International Journal of Distributed Informatique, 8(5–6):383–392, October/ December 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Inoue:1984:NTI

[142] Nobuo Inoue. New technologies for the Information Network System. Computer Networks: The International Journal of Distributed Informatique, 8(5–6):393– 412, October/December 1984. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Togawa:1984:IMS

[143] Akira Togawa. INS model system. Computer Networks: The International Journal of Distributed Informatique, 8 (5–6):413–420, October/December 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Mullender:1984:PRC

[144] Sape J. Mullender and Andrew S.

Tanenbaum. Protection and resource control in distributed operating systems. Computer Networks: The International Journal of Distributed Informatique, 8 (5–6):421–432, October/December 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Voydock:1984:SMT

[145] Victor L. Voydock and Stephen T. Kent.

Security mechanisms in a transport layer protocol. Computer Networks: The International Journal of Distributed Informatique, 8(5–6):433–449, October/ December 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Wegner:1984:SDF

[146] Lutz M. Wegner. Sorting a distributed file in a network. Computer Networks: The International Journal of Distributed Informatique, 8(5–6):451–461, October/ December 1984. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Sabnani:1984:VMS

[147] Krishnan Sabnani and Mischa Schwartz.

Verification of a multidestination selective repeat procedure. Computer Networks: The International Journal of Distributed Informatique, 8(5–6):463– 478, October/December 1984. CO-DEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Loveall:1980:LNF

[148] Robert M. Loveall and Richard D. Olson. Local networks in future military applications. In IEEE [207], pages 8–13. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Sauer:1980:LMN

[149] Anton M. Sauer, Heinz G. Schwaertzel, and Hans H. Bellm. Local microprocessor network for the office environment with an optical bus system. In IEEE [207], pages 17–20. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Schanning:1980:MPA

[150] Brian P. Schanning, Susan A. Powers, and John Kowalchuk. Memo: Privacy and authentication for the automated office. In IEEE [207], pages 21–30. CO-DEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Maglaris:1980:IBL

[151] Basil Maglaris and Tsvi Lissack. Integrated broadband local network architecture. In IEEE [207], pages 31–37. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Piney:1980:EIH

[152] C. Piney, C. Parkman, and F. Fluckiger. Endpoint interconnection of high-speed local packet networks in a virtual call environment. In IEEE [207], pages 41–50. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Kreager:1980:NLN

[153] Paul S. Kreager. No-modem local networks. In IEEE [207], pages 51–55. CO-DEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Nickens:1980:AHI

[154] Don O. Nickens, Thomas B. Genduso, and Stanley Y. W. Su. Architecture and hardware implementation of a prototype micronet. In IEEE [207], pages 56–64. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Dhas:1980:DRL

[155] C. Retna Dhas and David C. Rine. Design of a ring-based local area network for microcomputers. In IEEE [207], pages 65–68. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Warner:1980:CLN

[156] Clifford Warner. Connecting local networks to long haul networks: Issues in protocol design. In IEEE [207], pages 71–76. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Beauvais:1980:PRN

[157] Linda Beauvais and Dale H. Grit. Protocol for a ring network of heterogeneous computers. In IEEE [207], pages 77–85. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Horton:1980:MLC

[158] R. E. Horton and W. L. Miller. Multipoint link control protocol based on HDLC with source-destination addressing. In IEEE [207], pages 86–92. CO-DEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Watson:1980:SSC

[159] W. Bruce Watson. Simulation study of the configuration dependent performance of a prioritized, CSMA broadcast network. In IEEE [207], pages 95–103. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Cheng:1980:CSA

[160] Wing Kai Cheng and Geneva G. Belford. Clock synchronization algorithm for update synchronization on local broadcast networks. In IEEE [207], pages 104–110. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Wolfendale:1980:DDR

[161] G. L. Wolfendale. Development of a distributed resource sharing local computer network on CSIRONET. In IEEE [207], pages 111–119. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

Patton:1981:SGM

[162] Rex C. Patton. Support and growth of a manufacturing engineering coax computer network. In IEEE [208], pages 3–5. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Leech:1981:UMA

[163] Robert L. Leech and John R. Monastra. U. S. Military Academy Computer Local Area Network (CLAN). In IEEE [208], pages 6–9. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Czaplicki:1981:AAE

[164] Conrad S. Czaplicki. Advanced airborne executive. In IEEE [208], pages 10–12. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

deSilvaArroz:1981:ALC

[165] Guilherme de Silva Arroz. Arbiter for a local computer network with bus structure. In IEEE [208], pages 15–20. CO-DEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Li:1981:DAN

[166] Liang Li and Herman D. Hughes. Definition and analysis of a new shortest-delay protocol. In IEEE [208], pages 21–29. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Sidhu:1981:APG

[167] Deepinder P. Sidhu. Authentication protocols for general communication channels. In IEEE [208], pages 30–40. CO-DEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Ishizaka:1981:LLP

[168] Michihiro Ishizaka, Shigeo Nakatsuka, Takane Kakuno, and Tatsuki Ichihashi. Link level protocol and its firmware implementation in a ring network. In IEEE [208], pages 43–51. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Bergeron:1981:RMS

[169] Richard A. Bergeron, John L. Pokoski, and Robert D. Russell. RNS: a medium speed, low cost, local computer network. In IEEE [208], pages 52–58. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Liu:1981:ACP

[170] Then Tang Liu, Liang Li, and William R. Franta. Analysis of a conflict-free protocol based on node clusters. In IEEE [208], pages 61–72. CO-DEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Herskovitz:1981:NPT

[171] Jonas Herskovitz. Network protocol timing. In IEEE [208], pages 73–81. CO-DEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Katkin:1981:SCB

[172] Raphael D. Katkin and James G. Spring. Simulating a cable bus network in a multi-computer and large-scale application environment. In IEEE [208], pages 82–90. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Murphy:1981:CCM

[173] Jenanne L. Murphy. Centralized control and monitoring of a distributed local network. In IEEE [208], pages 93–99. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Graham:1981:BRL

[174] Martin Graham. Bit rates in linear common cable local networks. In IEEE [208], pages 100–101. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Husbands:1981:AVD

[175] Charles R. Husbands. Application of variable data rate transmission to local area fiber optic networks. In IEEE [208], pages 102–107. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

Trent:1982:WCP

[176] Barry A. Trent. Workstation clustering in a personal computer local network. In IEEE [209], pages 3–11. CO-DEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Gray:1982:TMT

[177] Alexander J. Gray and G. James Blaine III. Terranet: a modest terminalprocessor interconnect for the laboratory environment. In IEEE [209], pages 12–19. CODEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Sidhu:1982:LAN

[178] Deepinder P. Sidhu. Local area network design for military applications. In IEEE [209], pages 20–29. CODEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

OReilly:1982:DEF

[179] Peter J. P. O’Reilly, Joseph L. Hammond, Jay H. Schlag, and Dale N. Murray. Design of an emulation facility for performance studies of CSMA-type local networks. In IEEE [209], pages 35–43. CODEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Watson:1982:VDE

[180] W. Bruce Watson. Validation of a discrete event computer model of Network Systems Corporation’s Hyperchannel. In IEEE [209], pages 44–49. CO-DEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Ulm:1982:TTR

[181] John M. Ulm. Timed token ring local area network and its performance characteristics. In IEEE [209], pages 50–56. CODEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Wood:1982:LNA

[182] B. J. Wood, D. R. Thompson, L. D. Rogers, D. M. Bryant, and M. A. Malcolm. Local-area network architecture based on message-passing operating system concepts. In IEEE [209], pages 59–69. CODEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Moore:1982:DIL

[183] Lee C. Moore, Liudvikas Bukys, and James E. Heliotis. Design and implementation of a local network message passing protocol. In IEEE [209], pages 70–74. CODEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Gilligan:1982:GSA

[184] John M. Gilligan and John M. Vasak.

Generic security architecture for distributed systems. In IEEE [209], pages 75–82. CODEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Chafin:1982:LCN

[185] Roy L. Chafin. Local computer networks from a human factors perspective. In IEEE [209], pages 85–89. CO-DEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Protonotarios:1982:HPC

[186] E. N. Protonotarios, E. D. Sykas, and T. K. Apostolopoulos. Hybrid protocols for contention resolution in local area networks. In IEEE [209], pages 93–99. CODEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Marsan:1982:NMP

[187] M. Ajmone Marsan and D. Roffinella.

Nonpersistent M-CSMA protocols for multichannel local area networks. In IEEE [209], pages 100–109. CODEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Chu:1982:CBC

[188] Wesley W. Chu, Wilhelm Haller, and Kin K. Leung. Contention Based Channel Reservation (CBCR) protocol for high speed local networks. In IEEE [209], pages 110–117. CODEN CLCPDN. ISBN 0-89883-292-6. LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

Sharma:1983:SAC

[189] Roshan L. Sharma. Some architectural considerations for local area networks. In IEEE [210], pages 6–13. CODEN CLCPDN. ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE Service Cent. Piscataway, NJ, USA.

Murai:1983:KSU

[190] Jun Murai, Mario Tokoro, and Fumio Teraoka. KEIO S&Tnet: A UNIX campus network. In IEEE [210], pages 14–23. CODEN CLCPDN. ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE Service Cent. Piscataway, NJ, USA.

Minnich:1983:ETU

[191] N. M. Minnich and Charles J. Cotton. Evaluation of two Unibus Ethernet controllers. In IEEE [210], pages 29–36. CODEN CLCPDN. ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE Service Cent. Piscataway, NJ, USA.

Dahlberg:1983:PSF

[192] Arne Dahlberg. Performance study of fast loop networks. In IEEE [210], pages 37–43. CODEN CLCPDN. ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE Service Cent. Piscataway, NJ, USA.

Rosemberg:1983:PDC

[193] Flavia Rosemberg and Smil Ruhman.

Prioritized dialogue in CSMA/CD networks with provision for message cancellation. In IEEE [210], pages 47–54. CODEN CLCPDN. ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE Service Cent. Piscataway, NJ, USA.

Wah:1983:EPL

[194] Benjamin W. Wah and Jie-Yong Juang.

Efficient protocol for load balancing on CSMA/CD networks. In IEEE [210], pages 55–61. CODEN CLCPDN. ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE Service Cent. Piscataway, NJ, USA.

Maxemchuk:1983:SCM

[195] N. F. Maxemchuk. Some characteristics of movable slot TDM. In IEEE [210], pages 62–69. CODEN CLCPDN. ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE Service Cent. Piscataway, NJ, USA.

Sintonen:1983:TLU

[196] Leo Sintonen. Teletex-service for LAN users in a combined LAN-PDN interconnection environment. In IEEE [210], pages 75–82. CODEN CLCPDN.

ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE Service Cent. Piscataway, NJ, USA.

Lu:1983:CNA

[197] L. Y. Lu. Contextual naming, addressing and routing in interconnection. In IEEE [210], pages 83–87. CODEN CLCPDN. ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE Service Cent. Piscataway, NJ, USA.

Lenders:1984:LAN

[198] Patrick M. Lenders, Gearold R. Johnson, and Patrick A. Fitzhorn. Local area network for the College of Engineering at CSU. In IEEE [211], pages 5–9. CODEN CLCPDN. ISBN 0-8186-0589-8 (paperback), 0-8186-8589-1. LCCN TK 5105.7 C66 1984. IEEE Service Cent. Piscataway, NJ, USA.

Haugdahl:1984:KIS

[199] J. Scott Haugdahl. Key issues in selecting personal computer local networks. In IEEE [211], pages 10–14. CODEN CLCPDN. ISBN 0-8186-0589-8 (paperback), 0-8186-8589-1. LCCN TK 5105.7 C66 1984. IEEE Service Cent. Piscataway, NJ, USA.

Chlamtac:1984:SCI

[200] I. Chlamtac and O. Ganz. Study of communication interfaces for multiple bus networks. In IEEE [211], pages 18–25. CODEN CLCPDN. ISBN 0-8186-0589-8 (paperback), 0-8186-8589-1. LCCN TK 5105.7 C66 1984. IEEE Service Cent. Piscataway, NJ, USA.

Sethi:1984:PAT

[201] Adarshpal S. Sethi and Tuncay Saydam.

Performance analysis of token ring local area networks. In IEEE [211], pages 26–31. CODEN CLCPDN. ISBN 0-8186-0589-8 (paperback), 0-8186-8589-1. LCCN TK 5105.7 C66 1984. IEEE Service Cent. Piscataway, NJ, USA.

Jayasumana:1984:LTS

[202] Anura P. Jayasumana and P. David Fisher. Low-overhead token-passing scheme for bus networks. In IEEE [211], pages 32–38. CODEN CLCPDN. ISBN 0-8186-0589-8 (paperback), 0-8186-8589-1. LCCN TK 5105.7 C66 1984. IEEE Service Cent. Piscataway, NJ, USA.

Molloy:1984:CRC

[203] Michael K. Molloy. Collision resolution on the CSMA/CD bus. In IEEE [211], pages 44–47. CODEN CLCPDN. ISBN 0-8186-0589-8 (paperback), 0-8186-8589-1. LCCN TK 5105.7 C66 1984. IEEE Service Cent. Piscataway, NJ, USA.

Ibe:1984:HHB

[204] Oliver C. Ibe and Pin-Yee Chen. Hybrinet: a hybrid bus and token ring network. In IEEE [211], pages 48–54. CO-DEN CLCPDN. ISBN 0-8186-0589-8 (paperback), 0-8186-8589-1. LCCN TK 5105.7 C66 1984. IEEE Service Cent. Piscataway, NJ, USA.

Goel:1984:HFP

[205] R. K. Goel and A. K. El-Hakeem.

Hybrid FARA/CSMA-CD protocol for voice-data integration. In IEEE [211], pages 55–60. CODEN CLCPDN.

ISBN 0-8186-0589-8 (paperback), 0-8186-8589-1. LCCN TK 5105.7 C66 1984. IEEE Service Cent. Piscataway, NJ, USA.

IEEE:1979:CLC

[206] IEEE, editor. 4th Conference on Local Computer Networks, Marquette Inn, Minneapolis, Minnesota, October 22–23, 1979. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1979. ISBN ???? LCCN ????

IEEE:1980:CLC

[207] IEEE, editor. 5th Conference on Local Computer Networks, Marquette Inn, Minneapolis, Minnesota, October 6–7, 1980. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1980. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1980. IEEE catalog no. 80CH1542-0.

IEEE:1981:CLC

[208] IEEE, editor. 6th Conference on Local Computer Networks, Hilton Inn, Minneapolis, Minnesota, October 12–14, 1981. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. CODEN CLCPDN. ISBN ???? LCCN TK 5105.5 C66 1981. IEEE catalog no. 81CH1690-7.

IEEE:1982:CLC

[209] IEEE, editor. 7th Conference on Local Computer Networks: Hilton Inn, Minneapolis, Minnesota, October 11–13, 1982. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1982. CO-DEN CLCPDN. ISBN 0-89883-292-6.

LCCN TK 5105.5 C66 1982. IEEE catalog no. 82CH1800-2. Computer Society order no. 433.

IEEE:1983:CLC

[210] IEEE, editor. 8th Conference on Local Computer Networks, Hilton Inn, Minneapolis, Minnesota, October 17–19, 1983. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1983. CODEN CLCPDN. ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE catalog no. 83CH1931-5. Computer Society order no. 500.

IEEE:1984:CLC

[211] IEEE, editor. 9th Conference on Local Computer Networks, Minneapolis Plaza Hotel, Minneapolis, Minnesota, October 8–10, 1984. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. CODEN CLCPDN. ISBN 0-8186-0589-8 (paperback), 0-8186-8589-1. LCCN TK 5105.7 C66 1984. IEEE catalog no. 82CH2081-8. IEEE Computer Society order no. 589.